# Dynamics of the general $Q$-tensor model interacting with a rigid body

### v1

@@ 7. Local strong well-posedness for large data
$F_{3}(\widehat{z},\widehat{p})$ $\coloneqq-m_{\mathcal{S}}(\omega\times\ell)+\mu\ell+\int_{\partial\mathcal{S}_{0}}\bigl{(}(\mathrm{T}-\mathcal{T})(v,p,Q)-\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)\bigr{)}n\,\mathrm{d}\Gamma\enspace\text{and}\enspace$
$F_{4}(\widehat{z},\widehat{p})$ $\coloneqq\omega\times(J_{0}\omega)+\mu\omega+\int_{\partial\mathcal{S}_{0}}y\times\bigl{(}(\mathrm{T}-\mathcal{T})(v,p,Q)-\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)\bigr{)}n\,\mathrm{d}\Gamma.$
We now fix $T_{0}$, $R_{0}>0$ and consider $T\in(0,T_{0}]$ and $R\in(0,R_{0}]$. Recalling $\mathbb{E}_{T}$, $\mathbb{E}_{T}^{p}$ and $\tilde{\mathbb{E}}_{T}$ from (6.2) and (6.3), we introduce the set $\mathcal{K}_{T}^{R}$ together with the solution map $\Phi_{T}^{R}$ defined by
$\mathcal{K}_{T}^{R}\coloneqq\left\{(\tilde{z},\tilde{p})\in\prescript{}{0}{\mathbb{E}_{T}}\times\mathbb{E}_{T}^{p}:\|(\tilde{z},\tilde{p})\|_{\tilde{\mathbb{E}}_{T}}\leq R\right\}\enspace\text{and}\enspace\Phi_{T}^{R}\colon\mathcal{K}_{T}^{R}\to\prescript{}{0}{\mathbb{E}_{T}}\times\mathbb{E}_{T}^{p},\enspace\text{where}\enspace\Phi_{T}^{R}(\tilde{z},\tilde{p})\coloneqq(\widehat{z},\widehat{p}),$
and $(\widehat{z},\widehat{p})$ is the solution to (7.1) with $F_{1}(\tilde{z},\tilde{p})$, $F_{2}(\tilde{z})$, $F_{3}(\tilde{z},\tilde{p})$ and $F_{4}(\tilde{z},\tilde{p})$ for $(\tilde{z},\tilde{p})\in\mathcal{K}_{T}^{R}$. Thanks to the maximal $\mathrm{L}^{p}$-regularity from Theorem 6.1(a), $\Phi_{T}^{R}$ is well-defined if $(F_{1},F_{2},F_{3},F_{4})\in\mathbb{F}_{T}$, with $\mathbb{F}_{T}$ as made precise in (6.1). By construction, the existence and uniqueness of a solution to (5.11)–(5.13) is equivalent with $\Phi_{T}^{R}$ possessing a unique fixed point upon adding the reference solution $(z^{*},p^{*})$ from Section 7.
Next, we address the nonlinear estimates. For $T\in(0,T_{0}]$, $R\in(0,R_{0}]$, let $(\tilde{z},\tilde{p})\in\mathcal{K}_{T}^{R}$. With $(z^{*},p^{*})$ from Section 7, we set $(z,p)\coloneqq(\tilde{v}+v^{*},\tilde{Q}+Q^{*},\tilde{\ell}+\ell^{*},\tilde{\omega}+\omega^{*},\tilde{p}+p^{*})$. As observed in [9, Remark 5.2], for $C_{T}^{*}\coloneqq\|z^{*}\|_{\mathbb{E}_{T}}+\|p^{*}\|_{\mathbb{E}_{T}^{p}}$, it is valid that $C_{T}^{*}\to 0$ as $T\to 0$. By construction, it also follows that
(7.3) $\|(z,p)\|_{\tilde{\mathbb{E}}_{T}}\leq R+C_{T}^{*}.$
@@
In the sequel, we derive useful embeddings. The assertion of Section 7(a) is a consequence of the mixed derivative theorem, see e. g. [33, Corollary 4.5.10], and Sobolev embeddings. For part (b), we use [4, Theorem III.4.10.2] joint with interpolation theory and embeddings for the Besov spaces, see for example [40, Theorem 4.6.1].
[…]
@@
    $\mathbb{E}_{T}\hookrightarrow\mathrm{H}^{\theta,p}\bigl{(}0,T;\mathrm{H}^{2(1-\theta)}(\mathcal{F}_{0})^{3}\times\mathrm{H}^{2(1-\theta)+1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}\bigr{)}\enspace\text{and}\enspace\mathbb{E}_{T}^{Q}\hookrightarrow\mathrm{L}^{\infty}\bigl{(}0,T;\mathrm{H}^{2}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\bigr{)}.$
@@
In the sequel, we discuss useful estimates of $X$, $Y$ and related terms. For $(\ell_{i},\omega_{i})\in\mathrm{W}^{1,p}(0,T)^{6}$, the subscript $i$ is used to denote objects associated to $(\ell_{i},\omega_{i})$. Besides, we employ $\|\cdot\|_{\infty}$ and $\|\cdot\|_{\infty,\infty}$ for the $\mathrm{L}^{\infty}$-norm in time and time and space, respectively. The assertions of Section 7(a) and (b) follow upon modifying the arguments in [18, Section 6.1]. For the estimates of the second derivatives of $X$ and $Y$, note that the diffeomorphisms associated to $\ell=\omega=0$ are the identity, so their second derivatives vanish. Section 7(b) and Section 7 then imply $\|\ell_{i}\|_{\infty}+\|\omega_{i}\|_{\infty}\leq C(R+C_{0}+C_{T}^{*})$. For a proof of Section 7(c), we refer to [18, Lemma 6.5], while Section 7(d) can be deduced from (c).
[…]
@@
    $\|\partial_{j}\partial_{k}(Y_{i})_{l}\|_{\infty,\infty}$ $\leq CT(R+C_{0}+C_{T}^{*})\enspace\text{and}\enspace$ $\|\partial_{j}\partial_{k}(Y_{i})_{l}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}\leq CT(R+C_{0}+C_{T}^{*}).$
@@
    $\|\mathbb{O}_{1}-\mathbb{O}_{2}\|_{\infty}\leq CT\cdot\|M_{1}-M_{2}\|_{\infty}\leq CT\cdot\|\omega_{1}-\omega_{2}\|_{\infty}\enspace\text{and}\enspace\|\mathbb{O}_{i}\|_{\infty}+\|\mathbb{O}_{i}^{\top}\|_{\infty}\leq C.$
@@
Below, we collect estimates of some terms from (7.2). Here terms associated to $z_{i}$ are indicated by (i). The estimates of the first four terms in Section 7(a) can be deduced from [18, Lemma 6.6], while the last two terms in (a) can be handled similarly as in [9, (6.8) and Proposition 6.6] upon invoking the estimates in $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))$ from Section 7 and $\|\tilde{Q}\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{2}(\mathcal{F}_{0}))}\leq T^{\nicefrac{{1}}{{p}}}\cdot\|\tilde{Q}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{H}^{2}(\mathcal{F}_{0}))}\leq CT^{\nicefrac{{1}}{{p}}}\cdot\|\tilde{z}\|_{\mathbb{E}_{T}}$ for $\tilde{Q}\in\prescript{}{0}{\mathbb{E}_{T}^{Q}}$ by Section 7(a). For the estimates in Section 7(b) concerning the Cauchy stress tensor part, we refer to [18, Lemma 6.6], while the estimate of the remaining part can be derived from mapping properties of $\mathcal{J}$ defined below, the estimate $\|\mathbb{O}-\operatorname{Id}_{3}\|_{\infty}\leq CT$, following in turn from Section 7(b), and a suitable expansion of the difference.
[…]
Let $p>4$, recall $(z^{*},p^{*})$ from Section 7, and consider $(z,p)=(\tilde{z}+z^{*},\tilde{p}+p^{*})$ for $(\tilde{z},\tilde{p})\in\mathcal{K}_{T}^{R}$. Then there exists $C(R,T)>0$, with $C(R,T)\to 0$ as $R\to 0$ and $T\to 0$, such that
@@
    $\|(\mathcal{L}_{1}-\Delta)v\|_{\mathbb{F}_{T}^{v}}$, $\|\mathcal{M}v\|_{\mathbb{F}_{T}^{v}}$, $\|\mathcal{N}(v)\|_{\mathbb{F}_{T}^{v}}$, $\|(\mathcal{G}-\nabla)p\|_{\mathbb{F}_{T}^{v}}$, $\|(\mathcal{L}_{2}-\Delta)Q\|_{\mathbb{F}_{T}^{Q}}$, $\|(\partial_{t}Y\cdot\nabla)Q\|_{\mathbb{F}_{T}^{Q}}\leq C(R,T)$.
@@
    For $\varepsilon>0$, define $\mathcal{J}\colon\mathrm{H}^{\varepsilon+\nicefrac{{1}}{{2}}}(\mathcal{F}_{0})^{3\times 3}\to\mathbb{R}^{6}$ by $\mathcal{J}(h)\coloneqq(\int_{\partial\mathcal{S}_{0}}hn\,\mathrm{d}\Gamma,\int_{\partial\mathcal{S}_{0}}y\times hn\,\mathrm{d}\Gamma)^{\top}$. Then there is $C>0$ with $\|\mathcal{J}((\mathrm{T}-\mathcal{T})(v,p,Q))\|_{p}\leq CT\cdot\|z\|_{\mathbb{E}_{T}}$.
The lemma below on Lipschitz estimates of the non-transformed terms follows from [21, Lemma 6.3].
[…]
@@
[…]
Consider $p>4$, and for the reference solution $(z^{*},p^{*})$ from Section 7 and $(\tilde{z},\tilde{p})\in\mathcal{K}_{T}^{R}$, let $(z,p)=(\tilde{z}+z^{*},\tilde{p}+p^{*})$. Then there exists $C(R,T)>0$ such that $C(R,T)\to 0$ as $R\to 0$ and $T\to 0$, and with $\|\mathcal{B}_{\tau_{\mathrm{h}}}(\tilde{Q}+Q^{*})+\mathcal{B}_{\sigma}(\tilde{Q}+Q^{*})-\operatorname{div}S_{\xi}(Q_{0})(\Delta-\operatorname{Id}_{3})(\tilde{Q}+Q^{*})\|_{\mathbb{F}_{T}^{v}}\leq C(R,T)$.
@@
The next proposition asserts the self-map and Lipschitz estimates of $F_{1}$. It is a consequence of Section 7, Section 7 and the observation that the estimates of the transformed lower order terms $\mathcal{B}_{\tau_{\mathrm{l}}}$ are conceptually similar to those in Section 7. The Lipschitz estimates follow in the same way as the self-map estimate, see also [18, Section 6.1] or [9, Section 6.2] for examples of such Lipschitz estimates.
@@
Let $p>4$, and for the reference solution $(z^{*},p^{*})$ from Section 7 and $(\tilde{z},\tilde{p})$, $(\tilde{z}_{i},\tilde{p}_{i})\in\mathcal{K}_{T}^{R}$, $i=1,2$, consider $(z,p)=(\tilde{z}+z^{*},\tilde{p}+p^{*})$ and $(z_{i},p_{i})=(\tilde{z}_{i}+z^{*},\tilde{p}_{i}+p^{*})$. Besides, recall the $T$-independent constant $C_{\mathrm{MR}}>0$ from Theorem 6.1(a). Then there are $C_{1}(R,T)$, $L_{1}(R,T)>0$ such that $C_{1}(R,T)<\nicefrac{{R}}{{8C_{\mathrm{MR}}}}$ for $T>0$ sufficiently small and $L_{1}(R,T)\to 0$ as $R\to 0$ and $T\to 0$, and so that
$\|F_{1}(\tilde{z},\tilde{p})\|_{\mathbb{F}_{T}^{v}}\leq C_{1}(R,T)\enspace\text{and}\enspace\|F_{1}(\tilde{z}_{1},\tilde{p}_{1})-F_{1}(\tilde{z}_{2},\tilde{p}_{2})\|_{\mathbb{F}_{T}^{v}}\leq L_{1}(R,T)\cdot\|(\tilde{z}_{1},\tilde{p}_{1})-(\tilde{z}_{2},\tilde{p}_{2})\|_{\tilde{\mathbb{E}}_{T}}.$
@@
[…]
Let $p>4$, and for the reference solution $(z^{*},p^{*})$ from Section 7 and $(\tilde{z},\tilde{p})\in\mathcal{K}_{T}^{R}$, let $(z,p)=(\tilde{z}+z^{*},\tilde{p}+p^{*})$ and $(z_{i},p_{i})=(\tilde{z}_{i}+z^{*},\tilde{p}_{i}+p^{*})$. Then there is $C(R,T)>0$, with $C(R,T)\to 0$ as $R\to 0$ and $T\to 0$, such that $\|\mathcal{S}(\tilde{v}+v^{*},\tilde{Q}+Q^{*})+\tilde{S}_{\xi}(Q_{0})\nabla(\tilde{v}+v^{*})\|_{\mathbb{F}_{T}^{Q}}\leq C(R,T)$.
@@
As in the proof of Section 7, the idea here is to split the difference suitably. In fact, invoking that $-S(\nabla v,Q)=\tilde{S}_{\xi}(Q)\nabla v$ from Section 3, we write
@@
Consider $p>4$, and for the reference solution $(z^{*},p^{*})$ from Section 7 as well as $(\tilde{z},\tilde{p})$, $(\tilde{z}_{i},\tilde{p}_{i})\in\mathcal{K}_{T}^{R}$, $i=1,2$, let $(z,p)=(\tilde{z}+z^{*},\tilde{p}+p^{*})$ and $(z_{i},p_{i})=(\tilde{z}_{i}+z^{*},\tilde{p}_{i}+p^{*})$. Denote by $C_{\mathrm{MR}}>0$ the $T$-independent constant from Theorem 6.1(a). Then there exist $C_{2}(R,T)$, $L_{2}(R,T)>0$ with $C_{2}(R,T)<\nicefrac{{R}}{{8C_{\mathrm{MR}}}}$ for $T>0$ sufficiently small and $L_{2}(R,T)\to 0$ as $R\to 0$ and $T\to 0$, and with
$\|F_{2}(\tilde{z})\|_{\mathbb{F}_{T}^{Q}}\leq C_{2}(R,T)\enspace\text{and}\enspace\|F_{2}(\tilde{z}_{1})-F_{2}(\tilde{z}_{2})\|_{\mathbb{F}_{T}^{Q}}\leq L_{2}(R,T)\cdot\|\tilde{z}_{1}-\tilde{z}_{2}\|_{\mathbb{E}_{T}}.$
@@
In view of Section 7, Section 7 and the fact that the estimates of $\mu Q$ are simple, we still need to estimate $((\tilde{v}+v^{*})\cdot\nabla)(\tilde{Q}+Q^{*})$, $(\tilde{Q}+Q^{*})^{2}-\operatorname{tr}((\tilde{Q}+Q^{*})^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}}$ and $-\operatorname{tr}((\tilde{Q}+Q^{*})^{2})(\tilde{Q}+Q^{*})$.
@@
Let $p>4$, recall the reference solution $(z^{*},p^{*})$ from Section 7, and for $(\tilde{z},\tilde{p})$, $(\tilde{z}_{i},\tilde{p}_{i})\in\mathcal{K}_{T}^{R}$, $i=1,2$, set $(z,p)=(\tilde{z}+z^{*},\tilde{p}+p^{*})$, $(z_{i},p_{i})=(\tilde{z}_{i}+z^{*},\tilde{p}_{i}+p^{*})$. For $C_{\mathrm{MR}}>0$ from Theorem 6.1(a), there are $C_{3}(R,T)$, $C_{4}(R,T)>0$, with $C_{3}(R,T)$, $C_{4}(R,T)<\nicefrac{{R}}{{8C_{\mathrm{MR}}}}$ for $T>0$ small enough, and $L_{3}(R,T)$, $L_{4}(R,T)>0$, with $L_{3}(R,T)$, $L_{4}(R,T)\to 0$ as $R\to 0$, $T\to 0$, so that
$\|F_{3}(\tilde{z},\tilde{p})\|_{\mathbb{F}_{T}^{\ell}}$ $\leq C_{3}(R,T),\enspace\|F_{3}(\tilde{z}_{1},\tilde{p}_{1})-F_{3}(\tilde{z}_{2},\tilde{p}_{2})\|_{\mathbb{F}_{T}^{\ell}}\leq L_{3}(R,T)\cdot\|(\tilde{z}_{1},\tilde{p}_{1})-(\tilde{z}_{2},\tilde{p}_{2})\|_{\tilde{\mathbb{E}}_{T}}\enspace\text{as well as}\enspace$
$\|F_{4}(\tilde{z},\tilde{p})\|_{\mathbb{F}_{T}^{\omega}}$ $\leq C_{4}(R,T),\enspace\|F_{4}(\tilde{z}_{1},\tilde{p}_{1})-F_{4}(\tilde{z}_{2},\tilde{p}_{2})\|_{\mathbb{F}_{T}^{\omega}}\leq L_{4}(R,T)\cdot\|(\tilde{z}_{1},\tilde{p}_{1})-(\tilde{z}_{2},\tilde{p}_{2})\|_{\tilde{\mathbb{E}}_{T}}.$
@@
We start with the proof of Theorem 5.1, so let $(\tilde{z},\tilde{p})\in\mathcal{K}_{T}^{R}$. Theorem 6.1(a) applied to (7.1) together with Section 7, Section 7 and Section 7 results in
$\|\Phi_{T}^{R}(\tilde{z},\tilde{p})\|_{\tilde{\mathbb{E}}_{T}}\leq C_{\mathrm{MR}}\cdot\|(F_{1}(\tilde{z},\tilde{p}),F_{2}(\tilde{z}),F_{3}(\tilde{z},\tilde{p}),F_{4}(\tilde{z},\tilde{p}))\|_{\mathbb{F}_{T}}\leq C_{\mathrm{MR}}C(R,T).$

### v2

@@ 7. Local strong well-posedness for large data
[…]
We now fix $T_{0}$, $R_{0}>0$ and consider $T\in(0,T_{0}]$ and $R\in(0,R_{0}]$. Recalling $\mathbb{E}_{T}$, $\mathbb{E}_{T}^{\pi}$ and $\tilde{\mathbb{E}}_{T}$ from (6.2) and (6.3), we introduce the set $\mathcal{K}_{T}^{R}$ together with the solution map $\Phi_{T}^{R}$ defined by
[…]
Next, we address the nonlinear estimates. For $T\in(0,T_{0}]$, $R\in(0,R_{0}]$, let $(\tilde{z},\tilde{\pi})\in\mathcal{K}_{T}^{R}$. With $(z^{*},\pi^{*})$ from Section 7, we set $(z,\pi)\coloneqq(\tilde{v}+v^{*},\tilde{Q}+Q^{*},\tilde{\ell}+\ell^{*},\tilde{\omega}+\omega^{*},\tilde{\pi}+\pi^{*})$. As observed in [9, Remark 5.2], for $C_{T}^{*}\coloneqq\|z^{*}\|_{\mathbb{E}_{T}}+\|\pi^{*}\|_{\mathbb{E}_{T}^{\pi}}$, it is valid that $C_{T}^{*}\to 0$ as $T\to 0$. By construction, it also follows that
(7.3) $\|(z,\pi)\|_{\tilde{\mathbb{E}}_{T}}\leq R+C_{T}^{*}.$
@@
In the sequel, we derive useful embeddings. The assertion of Section 7(a) is a consequence of the mixed derivative theorem, see e. g. [35, Corollary 4.5.10], and Sobolev embeddings. For part (b), we use [4, Theorem III.4.10.2] joint with interpolation theory and embeddings for the Besov spaces, see for example [44, Theorem 4.6.1].
Lemma 7.3.
@@
    $\mathbb{E}_{T}\hookrightarrow\mathrm{H}^{\theta,p}\bigl(0,T;\mathrm{H}^{2(1-\theta)}(\mathcal{F}_{0})^{3}\times\mathrm{H}^{2(1-\theta)+1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}\bigr)\kern 5.0pt\text{and}\kern 5.0pt\mathbb{E}_{T}^{Q}\hookrightarrow\mathrm{L}^{\infty}\bigl(0,T;\mathrm{H}^{2}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\bigr).$
@@
In the sequel, we discuss useful estimates of $X$, $Y$ and related terms. For $(\ell_{i},\omega_{i})\in\mathrm{W}^{1,p}(0,T)^{6}$, the subscript $i$ is used to denote objects associated to $(\ell_{i},\omega_{i})$. Besides, we employ $\|\cdot\|_{\infty}$ and $\|\cdot\|_{\infty,\infty}$ for the $\mathrm{L}^{\infty}$-norm in time and time and space, respectively. The assertions of Section 7(a) and (b) follow upon modifying the arguments in [19, Section 6.1]. For the estimates of the second derivatives of $X$ and $Y$, note that the diffeomorphisms associated to $\ell=\omega=0$ are the identity, so their second derivatives vanish. Section 7(b) and Section 7 then imply $\|\ell_{i}\|_{\infty}+\|\omega_{i}\|_{\infty}\leq C(R+C_{0}+C_{T}^{*})$. For a proof of Section 7(c), we refer to [19, Lemma 6.5], while Section 7(d) can be deduced from (c).
Lemma 7.4.
@@
    $\|\partial_{j}\partial_{k}(Y_{i})_{l}\|_{\infty,\infty}$ $\leq CT(R+C_{0}+C_{T}^{*})\kern 5.0pt\text{and}\kern 5.0pt$ $\|\partial_{j}\partial_{k}(Y_{i})_{l}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}\leq CT(R+C_{0}+C_{T}^{*}).$
@@
    $\|\mathbb{O}_{1}-\mathbb{O}_{2}\|_{\infty}\leq CT\cdot\|M_{1}-M_{2}\|_{\infty}\leq CT\cdot\|\omega_{1}-\omega_{2}\|_{\infty}\kern 5.0pt\text{and}\kern 5.0pt\|\mathbb{O}_{i}\|_{\infty}+\|\mathbb{O}_{i}^{\top}\|_{\infty}\leq C.$
@@
Below, we collect estimates of some terms from (7.2). Here terms associated to $z_{i}$ are indicated by (i). The estimates of the first four terms in Section 7(a) can be deduced from [19, Lemma 6.6], while the last two terms in (a) can be handled similarly as in [9, (6.8) and Proposition 6.6] upon invoking the estimates in $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))$ from Section 7 and $\|\tilde{Q}\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{2}(\mathcal{F}_{0}))}\leq T^{\nicefrac{{1}}{{p}}}\cdot\|\tilde{Q}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{H}^{2}(\mathcal{F}_{0}))}\leq CT^{\nicefrac{{1}}{{p}}}\cdot\|\tilde{z}\|_{\mathbb{E}_{T}}$ for $\tilde{Q}\in\prescript{}{0}{\mathbb{E}_{T}^{Q}}$ by Section 7(a). For the estimates in Section 7(b) concerning the Cauchy stress tensor part, we refer to [19, Lemma 6.6], while the estimate of the remaining part can be derived from mapping properties of $\mathcal{J}$ defined below, the estimate $\|\mathbb{O}-\operatorname{Id}_{3}\|_{\infty}\leq CT$, following in turn from Section 7(b), and a suitable expansion of the difference.
Lemma 7.5.
Let $p>4$, recall $(z^{*},\pi^{*})$ from Section 7, and consider $(z,\pi)=(\tilde{z}+z^{*},\tilde{\pi}+\pi^{*})$ for $(\tilde{z},\tilde{\pi})\in\mathcal{K}_{T}^{R}$. Then there exists $C(R,T)>0$, with $C(R,T)\to 0$ as $R\to 0$ and $T\to 0$, such that
@@
    $\|(\mathcal{L}_{1}-\Delta)v\|_{\mathbb{F}_{T}^{v}}$, $\|\mathcal{M}v\|_{\mathbb{F}_{T}^{v}}$, $\|\mathcal{N}(v)\|_{\mathbb{F}_{T}^{v}}$, $\|(\mathcal{G}-\nabla)\pi\|_{\mathbb{F}_{T}^{v}}$, $\|(\mathcal{L}_{2}-\Delta)Q\|_{\mathbb{F}_{T}^{Q}}$, $\|(\partial_{t}Y\cdot\nabla)Q\|_{\mathbb{F}_{T}^{Q}}\leq C(R,T)$.
@@
    For $\varepsilon>0$, define $\mathcal{J}\colon\mathrm{H}^{\varepsilon+\nicefrac{{1}}{{2}}}(\mathcal{F}_{0})^{3\times 3}\to\mathbb{R}^{6}$ by $\mathcal{J}(h)\coloneqq(\int_{\partial\mathcal{S}_{0}}hn\,\mathrm{d}\Gamma,\int_{\partial\mathcal{S}_{0}}y\times hn\,\mathrm{d}\Gamma)^{\top}$. Then there is $C>0$ with $\|\mathcal{J}((\mathrm{T}-\mathcal{T})(v,\pi,Q))\|_{p}\leq CT\cdot\|z\|_{\mathbb{E}_{T}}$.
The lemma below on Lipschitz estimates of the non-transformed terms follows from [22, Lemma 6.3].
Lemma 7.6.
@@
Lemma 7.7.
Consider $p>4$, and for the reference solution $(z^{*},\pi^{*})$ from Section 7 and $(\tilde{z},\tilde{\pi})\in\mathcal{K}_{T}^{R}$, let $(z,\pi)=(\tilde{z}+z^{*},\tilde{\pi}+\pi^{*})$. Then there exists $C(R,T)>0$ such that $C(R,T)\to 0$ as $R\to 0$ and $T\to 0$, and with $\|\mathcal{B}_{\tau_{\mathrm{h}}}(\tilde{Q}+Q^{*})+\mathcal{B}_{\sigma}(\tilde{Q}+Q^{*})-\operatorname{div}S_{\xi}(Q_{0})(\Delta-\operatorname{Id}_{3})(\tilde{Q}+Q^{*})\|_{\mathbb{F}_{T}^{v}}\leq C(R,T)$.
@@
The next proposition asserts the self-map and Lipschitz estimates of $F_{1}$. It is a consequence of Section 7, Section 7 and the observation that the estimates of the transformed lower order terms $\mathcal{B}_{\tau_{\mathrm{l}}}$ are conceptually similar to those in Section 7. The Lipschitz estimates follow in the same way as the self-map estimate, see also [19, Section 6.1] or [9, Section 6.2] for examples of such Lipschitz estimates.
Proposition 7.8.
@@
$\|F_{1}(\tilde{z},\tilde{\pi})\|_{\mathbb{F}_{T}^{v}}\leq C_{1}(R,T)\kern 5.0pt\text{and}\kern 5.0pt\|F_{1}(\tilde{z}_{1},\tilde{\pi}_{1})-F_{1}(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\mathbb{F}_{T}^{v}}\leq L_{1}(R,T)\cdot\|(\tilde{z}_{1},\tilde{\pi}_{1})-(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\tilde{\mathbb{E}}_{T}}.$
@@
Lemma 7.9.
Let $p>4$, and for the reference solution $(z^{*},\pi^{*})$ from Section 7 and $(\tilde{z},\tilde{\pi})\in\mathcal{K}_{T}^{R}$, let $(z,\pi)=(\tilde{z}+z^{*},\tilde{\pi}+\pi^{*})$ and $(z_{i},\pi_{i})=(\tilde{z}_{i}+z^{*},\tilde{\pi}_{i}+\pi^{*})$. Then there is $C(R,T)>0$, with $C(R,T)\to 0$ as $R\to 0$ and $T\to 0$, such that $\|\mathcal{S}(\tilde{v}+v^{*},\tilde{Q}+Q^{*})+\tilde{S}_{\xi}(Q_{0})\nabla(\tilde{v}+v^{*})\|_{\mathbb{F}_{T}^{Q}}\leq C(R,T)$.
@@
As in the proof of Section 7, the idea here is to split the difference suitably. In fact, invoking that $S(\nabla v,Q)=\tilde{S}_{\xi}(Q)\nabla v$ from Section 3, we write
@@
$\|F_{2}(\tilde{z})\|_{\mathbb{F}_{T}^{Q}}\leq C_{2}(R,T)\kern 5.0pt\text{and}\kern 5.0pt\|F_{2}(\tilde{z}_{1})-F_{2}(\tilde{z}_{2})\|_{\mathbb{F}_{T}^{Q}}\leq L_{2}(R,T)\cdot\|\tilde{z}_{1}-\tilde{z}_{2}\|_{\mathbb{E}_{T}}.$
@@
In view of Section 7, Section 7 and the fact that the estimates of $\mu Q$ are simple, we still need to estimate $((\tilde{v}+v^{*})\cdot\nabla)(\tilde{Q}+Q^{*})$, $(\tilde{Q}+Q^{*})^{2}-\operatorname{tr}((\tilde{Q}+Q^{*})^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}}$ and $\operatorname{tr}((\tilde{Q}+Q^{*})^{2})(\tilde{Q}+Q^{*})$.
@@
$\|F_{3}(\tilde{z},\tilde{\pi})\|_{\mathbb{F}_{T}^{\ell}}$ $\leq C_{3}(R,T),\kern 5.0pt\|F_{3}(\tilde{z}_{1},\tilde{\pi}_{1})-F_{3}(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\mathbb{F}_{T}^{\ell}}\leq L_{3}(R,T)\cdot\|(\tilde{z}_{1},\tilde{\pi}_{1})-(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\tilde{\mathbb{E}}_{T}}\kern 5.0pt\text{as well as}\kern 5.0pt$
$\|F_{4}(\tilde{z},\tilde{\pi})\|_{\mathbb{F}_{T}^{\omega}}$ $\leq C_{4}(R,T),\kern 5.0pt\|F_{4}(\tilde{z}_{1},\tilde{\pi}_{1})-F_{4}(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\mathbb{F}_{T}^{\omega}}\leq L_{4}(R,T)\cdot\|(\tilde{z}_{1},\tilde{\pi}_{1})-(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\tilde{\mathbb{E}}_{T}}.$
@@
We start with the proof of Theorem 5.1, so let $(\tilde{z},\tilde{\pi})\in\mathcal{K}_{T}^{R}$. Theorem 6.1(a) applied to (7.1) together with Section 7, Section 7 and Section 7 results in
$\|\Phi_{T}^{R}(\tilde{z},\tilde{\pi})\|_{\tilde{\mathbb{E}}_{T}}\leq C_{\mathrm{MR}}\cdot\|(F_{1}(\tilde{z},\tilde{\pi}),F_{2}(\tilde{z}),F_{3}(\tilde{z},\tilde{\pi}),F_{4}(\tilde{z},\tilde{\pi}))\|_{\mathbb{F}_{T}}\leq C_{\mathrm{MR}}C(R,T).$